OpenStudy (anonymous):

@Hero The sides of a triangle have lengths of 8 cm, 15 cm, and 17 cm. Which of the following statements is true? A. The triangle is a right triangle, because 8 + 15 > 17. <---- B. The triangle is a right triangle, because 8^2 + 15^2 = 17^2. C. The triangle is a right triangle, because (8 + 15)^2 = 17^2. D. The triangle is not a right triangle. A is the only one that makes sense to me. :)

hero (hero):

Why does that one make sense to you? What reasoning did you use to arrive at your answer choice?

OpenStudy (solomonzelman):

\[8^2+15^2=17^2\]

OpenStudy (anonymous):

Hold on, there's a snake outside my door!

OpenStudy (solomonzelman):

You live in India ?

hero (hero):

In general, If a^2 + b^2 = c^2, then you have a right triangle if a^2 + b^2 > c^2 then you have an acute triangle if a^2 + b^2 < c^2 then you have an obtuse triangle

OpenStudy (anonymous):

No, the us. My grandmother is freaking out because it's on our screen dorr and we looked out the window and bam! its there sao im trying to find out what kind it is and whethr or not its pisness.

hero (hero):

*poisonous

OpenStudy (solomonzelman):

probably not :)

hero (hero):

I would never assume a snake or spider isn't poisonous. I don't trust either one of 'em.

OpenStudy (solomonzelman):

true

OpenStudy (anonymous):

I'm almost positive it's this one... http://en.wikipedia.org/wiki/Seminatrix :) Now on with my explanations...

OpenStudy (anonymous):

The sides of a triangle have lengths of 8 cm, 15 cm, and 17 cm. Which of the following statements is true? A. The triangle is a right triangle, because 8 + 15 > 17. B. The triangle is a right triangle, because 8^2 + 15^2 = 17^2. C. The triangle is a right triangle, because (8 + 15)^2 = 17^2. D. The triangle is not a right triangle. A ~ 8+15=23>17 So TRUE! :) B ~ (8^2=64) + (15^2=225) = 289 17^2= 289 Hm..... That one is true.... C ~ (15+8=23)^2=529 17^2-289 SO...FALSE! :) @Hero

OpenStudy (anonymous):

@Hero

hero (hero):

I already gave you parameters above: If a^2 + b^2 = c^2, then you have a right triangle if a^2 + b^2 > c^2 then you have an acute triangle if a^2 + b^2 < c^2 then you have an obtuse triangle But there's another parameter to consider for option D a + b > c b + c > a a + c > b

hero (hero):

The first set of parameters are for determining what kind of triangle you have. The second set determines if you have a triangle or not.

hero (hero):

For a + b > c b + c > a a + c > b All three must be true in order for the triangle to exist.

OpenStudy (anonymous):

I believe it's B, after looking over what you wrote. :) @Hero